•The pressure of a gas in a tank is 3.2 atm at 22°C. If temp rises to 60°C, what will the pressure be?

•A 5.00-L air sample has a pressure of 107 kPa at a temperature of -50.0°C. If the temperature is raised to 102°C and the volume expands to 7.00 L, what will the new pressure be?

•Helium at 23°C with a volume of 5.60 L at 2.45 atm is changed to a pressure of 8.75 atm and cooled to 15°C. What is the volume?
Answer #1

2. P = Ti P2 Tz Here P, = 3.2aton, T=22+273=295k T2 = 60+273 = 333K, P2 = PIT2 = 3.28 333 =3.612 aten T 295 PUI = P22 Ti Tа.